Show that the maximum tension in the string of a simple pendulum, when the

amplitude θ is small, is mg(1+(θ)^2). At what position of the pendulum is the tension a

maximum?


1
Expert's answer
2021-05-09T13:04:40-0400


Tension of the string when the string reaches the mid point

"T=T_m"

"T_m=mg+\dfrac{mv^2}{l}"


Applying energy balance

"\dfrac{1}{2}mv^2=mgl(1-cos\theta)"

"v^2=2gl(1-cos\theta)"

"v^2=2gl\Big(2sin^2\dfrac{\theta}{2}\Big)"


Substituting the value of "v^2" in above equation

"T_m=mg+2mgl\dfrac{\Big(2sin^2\dfrac{\theta}{2}\Big)}{l}"

"T_m=mg+4mg\Big(sin^2\dfrac{\theta}{2}\Big)"

"T_m=mg\Big(1+4sin^2\dfrac{\theta}{2}\Big)"

for small "\theta, \space sin\theta\approx\theta"

"T_m=mg\Big(1+4\Big(\dfrac{\theta}{2}\Big)^2\Big)"

"T_m=mg(1+\theta^2)"


The tension of the string is maximum at extreme positions